Mr. Johnson rode his bike from his house to a store in Woodland Park at a rate of 10.3 miles per hour. He immediatly turned around and came home at an average rate of 22.5 miles per hour. If the total trip took him approximatley 5.15 hours, how far is it from his house to the store?
Answer by stanbon(48558)

Mr. Johnson rode his bike from his house to a store in Woodland Park at a rate of 10.3 miles per hour. He immediatly turned around and came home at an average rate of 22.5 miles per hour. If the total trip took him approximatley 5.15 hours, how far is it from his house to the store?
---------------------------------
House to Store DATA:
rate = 10.3 mph ; distance = x miles ; time = d/r = x/10.3
----------------------
Store to House DATA:
rate = 22.5 mph ; distance = x miles ; time = d/r = x/22.5
---------------------
Equation:
time + time = 5.15 hrs
x/10.3 + x/22.5 = 5.15
(22.5x + 10.3x) = 10.3*22.5*5.15
32.8x = 1193.5125
x = 36.39 miles (distance between house and store)
=========================
Cheers,
Stan H.